OpenStudy (anonymous):

Probability Question. How many ways can seven people A, B, C, D, E, F, and G sit on a bench if A wants to sit next to B and F wants to sit on one of the ends?

OpenStudy (anonymous):

I set up a possibliities chart and came up with 2*6*1*4*3*2*1 and got 288 ways. It's wrong though.

OpenStudy (raden):

48 ?

OpenStudy (anonymous):

It's supposed to be 480, but thanks anyways.

OpenStudy (raden):

ah yes, it is 480 case I (F on the right end) : AB____F = 4! 2! _AB___F = 4! 2! __AB__F = 4! 2! ___AB_F = 4! 2! ____ABF = 4! 2! TOTAL = 5 * 4! * 2! = 240 case II (F on the left side) FAB____ = 4! 2! F_AB___ = 4! 2! F__AB__ = 4! 2! F___AB_ = 4! 2! F____AB = 4! 2! TOTAL = 5 * 4! 2! = 240 240 + 240 = 480 ways

OpenStudy (anonymous):

thank you

OpenStudy (raden):

you're welcome